OpenStudy (logo):

Help pls? Will award medal! What is the solution to x2 + 3x + 25 = 0?

sam (.sam.):

Do you think the equation is factorable?

OpenStudy (angel.candy):

I'm not sure if you meant 2x instead of x2.

OpenStudy (logo):

No.. and its x^2

sam (.sam.):

x2 means \(x^2\)

OpenStudy (angel.candy):

Ah okay, thats what I originally thought, but I felt it was more basic. :P Sorry hehe

OpenStudy (joseiscoolya):

use wolframalpha.com and type in factor "your equation"

sam (.sam.):

The solution is actually imaginary :)

sam (.sam.):

Or no real solutions

OpenStudy (logo):

So I'll have to use the -b+- sqrt equation right?

sam (.sam.):

Yep, this to be exact \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

sam (.sam.):

Where a=1, b=3, c=25

OpenStudy (logo):

y = -b+- sqrt_/ b^2 - 4ac/2a y = -3 +- sqrt_/3^2 - 4*1*25 / 2*1 y = -3 +- sqrt_/ 9 - 4*25 / 2 y = -3 +- sqrt_/ 9-100 / 2 y = -3 +- sqrt/ -46 so like this?

OpenStudy (welshfella):

-3 +/- sqrt (9 - 4*25) = -3 +/- sqrt (-91)

OpenStudy (welshfella):

- oh all that divided by 2 !!!

OpenStudy (welshfella):

are you familiar with imaginary numbers? sqrt(-1) = i ?

OpenStudy (logo):

sooo I find the square root of -91 and then divide it by 2?

OpenStudy (welshfella):

you divide it all by 2 -3 +/- sqrt91 i ------------ 2

OpenStudy (logo):

oh ok

OpenStudy (welshfella):

i is the symbol for sqrt (-1)

OpenStudy (welshfella):

the roots consist of 2 complex numbers a complex number has a real and an imaginary part

OpenStudy (welshfella):

we can write them as